OpenStudy (anonymous):

Miguel is making an obstacle course for field day. At the end of every sixth of the course, there is a tire. At the end of every third of the course there is a cone. At the end of every half there is a hurdle. At which locations of the course will people need to go through more than obstacle? My daughter's homework

OpenStudy (barreraa):

The answers are: 2/6, 3/6, 4/6, and 6/6.

OpenStudy (barreraa):

@islandgurl

OpenStudy (anonymous):

explain ansr please

OpenStudy (barreraa):

We have three fractions with different denominators: sixths, thirds, and halves. The first step is to make all the denominators equal. In this case we want sixths. 1/3 = 2/6, and 1/2 = 3/6. Now we can start solving. 1. There are six tires at the following: 1/6, 2/6, 3/6, 4/6, 5/6, and 6/6. 2. There are three cones at the following (G.C.F.): 2/6 (or 1/3), 4/6 (or 2/3), and 6/6 (or 3/3). 3. There are two hurdles at the following (G.C.F.): 3/6 (or 1/2) and 6/6 (or 2/2). We simply look for common numbers. 1. At 2/6, there are two obstacles: a tire and a cone. 2. At 3/6, there are two obstacles: a tire and a hurdle. 3. At 4/6, there are two obstacles: a tire and a cone. 4. At 6/6, there are three obstacles: a tire, cone, and a hurdle. (hard!) The answers are: 2/6, 3/6, 4/6, and 6/6.

OpenStudy (anonymous):

thank you

OpenStudy (barreraa):

np :) :)

OpenStudy (barreraa):

http://brainly.com/question/195266